# Bio-inspired reconfigurable stereo vision for robotics using omnidirectional cameras

### v1

@@ IV Demonstration of the reconfigurable vision
To further demonstrate the potential application of our reconfigurable vision system, we deployed it on a metamorphic quadruped robot, Origaker, which is a reconfigurable robot capable of altering its shape and mobility features by shifting its closed eight-bar linkage trunk[24]. Two omnidirectional cameras are installed on the robot’s trunk linkage. Consequently, the relative position of the two omnidirectional cameras changes as the robot’s form transforms, acquiring different vision features. The three configurations (gecko, spider, and stick-bug), of Origaker that can carry the camera system are shown in Fig. 8 and described below.
@@
The gecko configuration provides a herbivorous-like spatial view, with a stereoscopic FOV of $\sim$76° and a monocular FOV of $\sim$316° (the widest one). The spider configuration provides a carnivorous-like spatial view, with a stereoscopic FOV of $\sim$136° and a monocular FOV of $\sim$256°. The stick-bug configuration provides a moderate spatial view, with a stereoscopic FOV of $\sim$150° (the widest one) and a monocular FOV of $\sim$242°. By adapting to these different configurations, the vision system on Origaker can exhibit varying perceptual characteristics, showcasing its versatility and robustness in different operational scenarios.
@@
Here, to exhibit the adaptability of the reconfigurable vision, the Origaker was placed in a complex scenario for approaching a specific object in a narrow corridor and inspecting it in detail. The robot started in the gecko configuration, utilizing its wide FOV for extensive sensing and searching (Fig. 8(a)). With a relatively limited blind zone, the robot can obscurely avoid obstacles in almost all directions. The narrow field angle sensing in the center allows for minimal stereoscopic environment sensing to assist in-depth perception. After detecting the target object, the Origaker switched to the spider configuration to gain peripheral depth sensing (Fig. 8(b)). The blind zone is still less than 180°, ensuring a broad range of sensing capabilities. This configuration helps the robot accurately measure distances to the corridor walls and avoid collisions. Upon approaching the target object, the Origaker switched to the stick-bug configuration for detailed close-range observation (Fig. 8(c)). This mode offers improved camera alignment, allowing for the matching of homologous points in the immediate vicinity, which is difficult for other configurations to achieve. However, the matching process is more challenging due to the high view angle difference, and there is a higher likelihood of distant objects being blocked by closer ones. The stick-bug configuration enables precise depth measurement of nearby objects, providing a detailed view for accurate shape detection.
@@ V Conclusions and future work
This research introduces a bio-inspired reconfigurable stereo vision system for robotics, using omnidirectional cameras to achieve adaptable visual features. Emulating animal visual adaptability, it overcomes the limitations of traditional fixed alignment of cameras and narrow FOV. Key innovations include a reconfigurable stereo vision strategy, a robust depth measurement system using a modified nonrectified geometrical method and deep neural network, and the successful deployment of the metamorphic robot Origaker. The system offers flexibility with different vision modes, such as a wide field for target seeking and a broad binocular field for detailed inspection, highlighting the potential of bio-inspired designs to enhance robotic vision adaptability in various scenarios. Future work will focus on developing a mechanical design that closely mimics the eye alignment mechanisms to deeply replicate the various visual strategies of different species. Combining embodied intelligence with this system will further improve the adaptability of our vision system.

### v2

@@ IV Demonstration of the reconfigurable vision
To further demonstrate the potential application of our reconfigurable vision system, we deployed it on a metamorphic quadruped robot, Origaker, which is a reconfigurable robot capable of altering its shape and mobility features by shifting its closed eight-bar linkage trunk[24]. Two omnidirectional cameras are installed on the robot’s trunk linkage. The relative pose relationship between the robot’s geometric model and cameras is demonstrated in Fig. 10. Consequently, the relative position of the two omnidirectional cameras changes as the robot’s form transforms, acquiring different vision features. The transitions between the visual stages are automatically executed upon receiving a human-issued control command,
Figure 10: Sketch and coordinate system of the metamorphic trunk of Origaker with two installed cameras.
Figure 11: Demonstration of the proposed reconfigurable vision system on the metamorphic robot Origaker. Measured depth value colored according to the color bar.In the figure, the head link configuration of Origaker and the vision system is shown, along with the actual scenarios and corresponding vision information acquired from (a) The Gecko form. (b) The Spider form. (c) The Stick-bug form.
Origaker’s torso employs a spatial closed-loop eight-bar linkage mechanism, exhibiting remarkable reconfigurability for mimicking structures of different creatures. This work focuses on three primary configurations: gecko, spider, and stick-bug modes. These configurations leverage the six-bar planar mechanism derived from the eight-bar spatial structure through kinematic degeneration (joints 7 and 8 are fixed). A modified Denavit-Hartenberg parameter method is adopted for robot modeling, where closed-loop equations are typically applied to solve joint angles for such mechanisms.
The closed-loop equation could be written as:
$T_{1}^{2}T_{2}^{3}T_{3}^{4}T_{4}^{5}T_{5}^{6}T_{6}^{1}=I$ (6)
where $T_{i}^{j}$ represents the homogeneous transformation matrix from joint $j$ to joint $i$, ensuring the loop closure of the mechanism. This closed-loop equation establishes a system of nonlinear equations, which is solved using MATLAB’s nonlinear least squares solver lsqnonlin under given $q_{i}$ conditions to determine all joint angles. Subsequent rigid transformations yield camera poses $\{C_{1}\}$ and $\{C_{2}\}$ in our reconfigurable vision system, as shown in Fig. 10.
@@
The gecko configuration provides a herbivorous-like spatial view, with a stereoscopic FOV of $\sim$76° and a monocular FOV of $\sim$207° (the widest one). The spider configuration provides a carnivorous-like spatial view, with a stereoscopic FOV of $\sim$136° and a monocular FOV of $\sim$87°. The stick-bug configuration provides a moderate spatial view, with a stereoscopic FOV of $\sim$140° (the widest one) and a monocular FOV of $\sim$95°. The FOV value mentioned here is calculated with actual robot setting parameters and an equivalent object distance $R$ = 3000 mm. By adapting to these different configurations, the vision system on Origaker can exhibit varying perceptual characteristics, showcasing its versatility and robustness in different operational scenarios.
@@
Here, to exhibit the adaptability of the reconfigurable vision, the Origaker was placed in a complex scenario to approach a specific object in a narrow corridor and inspect it in detail. The robot started in the gecko configuration, utilizing its wide FOV for extensive sensing and searching (Fig. 11(a)). With a relatively limited blind zone, the robot can obscurely avoid obstacles in almost all directions. The narrow field angle sensing in the center allows for minimal stereoscopic environment sensing to assist in-depth perception. After detecting the target object, the Origaker switched to the spider configuration to gain peripheral depth sensing (Fig. 11(b)). The blind zone is still less than 180°, ensuring a broad range of sensing capabilities. This configuration helps the robot accurately measure distances to the corridor walls and avoid collisions. Upon approaching the target object, the Origaker switched to the stick-bug configuration for detailed close-range observation (Fig. 11(c)). This mode optimizes camera alignment, facilitating the accurate matching of nearby homologous points, which other configurations struggle to achieve. However, the matching process becomes more challenging due to the significant difference in viewing angles, increasing the likelihood of occlusion, where distant objects are obstructed by closer ones. The stick-bug configuration, however, enhances depth measurement accuracy for nearby objects, offering a detailed perspective for precise shape reconstruction. Additionally, the transformation process and stereo vision features of each configuration are also demonstrated in Video 1.
@@ V Conclusions and future work
This research introduces a bio-inspired reconfigurable stereo vision system for robotics, using omnidirectional cameras to achieve adaptable visual features. Emulating animal visual adaptability, it overcomes the limitations of the traditional fixed alignment of cameras and narrow FOV. Key innovations include a reconfigurable stereo vision strategy, a robust depth measurement system using a modified nonrectified geometrical method and deep neural network, and the successful deployment of the metamorphic robot, Origaker. The system offers flexibility with different vision modes, such as a wide field for target seeking and a broad binocular field for detailed inspection, highlighting the potential of bio-inspired designs to enhance robotic vision adaptability in various scenarios. Future research will focus on refining the mechanical design to closely mimic eye alignment mechanisms, enabling a more comprehensive replication of the diverse visual strategies observed in different species. Additionally, having the system’s transformation further automated with the task requirement is also a promising direction.